Find a polynomial function in standard form with the following roots x= -2, 1(multiplicity 2), 0. Thanks
Answer by khai(18)

(x+2)(x-1)(x-1)x
= (x+2)(x^2-2x+1)x
= (x^3-2x^2+x+2x^2-4x+2)x
= (x^3-3x+2)x
= (x^4-3x^2+2x)